OpenStudy (anonymous):

Will Fan and Medal!!!!! Choose the correct simplification of the expression 6 times b all over a to the power of negative 9.

OpenStudy (anonymous):

a to the 9th power times b all over 6 6 times b all over a to the 9th power Already simplified 6a9b

OpenStudy (anonymous):

@Abhisar @Michele_Laino @CGGURUMANJUNATH @coolman500 @PizzaLover123

OpenStudy (anonymous):

@TheSmartOne @tkhunny @sammixboo @sleepyjess @ikram002p @shifuyanli

OpenStudy (anonymous):

@♪Chibiterasu @Smita12

TheSmartOne (thesmartone):

Translate this from words into an equation: 6 times b all over a to the power of negative 9. Hint: times means multiplication all over means there is fraction to the power of means \(\sf a^b\) and also \(\sf\Large x^{-y}=\frac{1}{x^y}\) so then: \(\sf\Large\frac{1}{a^{-b}}=a^b\)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so C @TheSmartOne

TheSmartOne (thesmartone):

Nope, do some work or rather show some work on how you arrived to that.

OpenStudy (melissa_something):

6 · b all over (fraction) a^ -9

OpenStudy (anonymous):

thanks*

OpenStudy (anonymous):

could u help on one more i will fan

OpenStudy (anonymous):

@Melissa_Something

OpenStudy (melissa_something):

give me a medal so i know it's real ;D

OpenStudy (anonymous):

i did

OpenStudy (melissa_something):

show me the problem,

OpenStudy (anonymous):

Choose the correct simplification of the expression -3x2(5x - 4x2 - 6). -7x4 + 2x3 - 9x2 12x4 - 15x3 + 18x2 12x4 + 15x3 + 18x2 -12x4 + 15x3 - 18x2

OpenStudy (melissa_something):

First you will apply the distributive property, simplify, and then combine like terms. You should get b

OpenStudy (anonymous):

thanks :)

TheSmartOne (thesmartone):

Her answer to the first question was wrong fyi.

TheSmartOne (thesmartone):

And giving and taking direct answers is against the code of conduct. http://openstudy.com/code-of-conduct